A quality control agent tested 8 pints of ice cream during the past hour. She found that one of the pints was not acceptable. Using the relative frequency probability method, find the probability that the next pint tested will be acceptable.

  1. 0.875
  2. 0.125
  3. 0.25
  4. 0.50
Answer #1

Total number of pints = 8

Number of acceptable pints = 8 - 1 = 7

P(next pint tested will be acceptable) = Number of acceptable pints / Total number of pints tested

= 7/8

= 0.875